OpenStudy (anonymous):

Jennifer's basketball team played 18 games this season, and they won twice as many games as they lost. How many games did they lose?

OpenStudy (anonymous):

9?I think

OpenStudy (anonymous):

no wait lol

OpenStudy (anonymous):

6

OpenStudy (anonymous):

18=l+w w=2*l therefore 18=l+2*l solve for l

OpenStudy (anonymous):

6